Write an equation in slope-intercept form for the line that passes through the points (6, -3) and (0, 2).
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Write an equation in slope-intercept form for the line that passes through the points (6, -3) and (0, 2).
===============================
Find the slope, m
m = (y change)/(x change)
----
Then, y-y1 = m*(x-x1) where (x1,y1) is either point.
==============================
Know this like you know 2+2 = 4

Answer by ikleyn(52794)

.

There is the lesson in this site, which explains how to solve such problems

    - Equation for a straight line in a coordinate plane passing through two given points